OpenStudy (anonymous):

evaluate. 2 integrals

hartnn (hartnn):

did you try \( u= \sqrt y \)

OpenStudy (anonymous):

the u sub method

hartnn (hartnn):

for 1st one, try \(u =\sqrt y \) and for 2nd one, try \(u =\sqrt x \)

OpenStudy (anonymous):

for the first one, my du is 1/2 y ^-1/2 dy. and for the 2nd, du = 1/2 x ^-1/2

hartnn (hartnn):

so 2du = dy/ sqrt y right ?

hartnn (hartnn):